Kim wants to determine a 80 percent confidence interval for the true proportion of high school students in the area who attend their home basketball games. How large of a sample must she have to get a margin of error less than 0.03? [To find n, use the value p* = 1/2 for the sample proportion and the values for z* from a z-table or t-table.]

[Round to the smallest integer that works.] n =
Answer #1

solution:-

given that proportion p = 1/2 = 0.5

80% confidence for z is 1.28

margin of error = 0.03

formula for sample n

n = p * q * (Z/E)^2

= 0.5 * 0.5 * (1.28/0.03)^2
= 455